An automobile starting from the rest speeds up to 15 m/sec. with a constant acceleration of 1.2 m/sec^2 , runs at fast speed for sometime and finally comes to the rest with the deceleration of 1.5 m/sec ^2 . If the total distance travelled is 400 m find the total time required. *

1
Expert's answer
2022-02-14T14:13:54-0500

Solution

For first case

Required time

"t_1=\frac{v}{a}\\=\frac{15}{1.2}\\=12.5sec"

Now for second case

"t_2=\frac{v}{a}\\=\frac{15}{1.5}\\=10sec"

So total time required

"t=t_1+t_2\\=12.5+10=22.5sec"